OpenStudy (anonymous):

I need help with this problem: /-7-8/+5 The lines are absolute value signs. Is it 20 or 4?

OpenStudy (anonymous):

what does it equal?

OpenStudy (anonymous):

=0?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

|-7-8|+5 = |-15| + 5 = 15+5 = 20

OpenStudy (anonymous):

@Denebel we dont know what it equals...

OpenStudy (anonymous):

Mrbonez, he was asking what it equals. And it's 20.